OpenStudy (anonymous):

Triangle ABC has side lengths 3, 4, and 5. Do the side lengths form a Pythagorean triple? Explain.

OpenStudy (smarty):

Yes they do, because when you plug in the numbers for A,B, and C, it satisfies the equation.

OpenStudy (anonymous):

yes because a pythagorean triple means that if you use a^2 + b^2 = c^2 the equation works. 3 squared (9) + 4 squared (16) = 5 squared (25) .

OpenStudy (anonymous):

A. No, they do not form a Pythagorean triple; although \[3^{2}+4^{2}=5^{2}\], the side lengths do not meet the other requirements of a Pythagorean triple. B. No, they do not form a Pythagorean triple;\[3^{2 + 4^{2 \neq 5^{2} }}\] . C. Yes, they form a Pythagorean triple\[3^{2}+4^{2}=5^{2}\] ; and 3, 4, and 5 are all nonzero whole numbers.

OpenStudy (anonymous):

and the b option came out wrong its suppose to look like the A equation but with the line throught the = sign.

OpenStudy (smarty):

C is your answer then...

OpenStudy (anonymous):

okay thank you so much :)